OpenStudy (anonymous):

the 20th term of the sequence √5,√20,√45, ... is ?

OpenStudy (amistre64):

sqrt(5,+15=20,+25=45,+35=80,+45=125) maybe?

OpenStudy (amistre64):

it consistent, but is it right?

OpenStudy (anonymous):

i dont know .. i have options 20 √5 , 10V5

OpenStudy (amistre64):

20 sqrt(5) seems the best bet

OpenStudy (anonymous):

i think 1st is correct....but i dont know to find complex no. sequence..

OpenStudy (amistre64):

1*sqrt(5) 2* " 3* " 4* " ....

OpenStudy (anonymous):

20 sqrt 5 is correct

OpenStudy (amistre64):

16 5 ---- 80 lol ... i wonder if mine works :)

OpenStudy (anonymous):

how to do it ? pls tell me..

OpenStudy (amistre64):

you find a pattern, and see which options fits

OpenStudy (anonymous):

the nth term is n * sqrt5

OpenStudy (anonymous):

okie...

OpenStudy (anonymous):

so 20√5..

OpenStudy (anonymous):

right

OpenStudy (anonymous):

then why the 2nd term is not 2√5 ??

OpenStudy (anonymous):

is it just multiplying 1st term with nth term?

OpenStudy (anonymous):

it is sqrt20 = sqrt4 *sqrt5 = 2 sqrt 5

OpenStudy (anonymous):

okie ... got it ...thanks a lot.. :)

OpenStudy (anonymous):

third term = 3 * sqrt5 = sqrt9 * sqrt5 = sqrt 45

OpenStudy (anonymous):

you r welcome